PDS_VERSION          = PDS3
RECORD_TYPE          = STREAM
OBJECT               = TEXT
PUBLICATION_DATE       = 1998-05-07
NOTE                   = "MGS RST Instrument Health Report HEA8127A.TXT"
END_OBJECT           = TEXT
END


Miscellaneous
=============

     The amount of spacecraft engineering data collected from yesterday was
very small (about 250 kB, as opposed to 50 MB in other recent days).  We
assumed this is because of the approaching solar conjunction.


Engineering Data
================

Engineering Data Statistical Summary from CSV2STAT:
 First SCET: 1998-05-06T06:48:48.891
 Last SCET:  1998-05-06T23:34:36.385

CHANNL CHANNEL TITLE  DN-LO DN-HI   EU-LO     EU-HI      EU-AVG/STDEV     RECDS
------ -------------  ----- ----- --------- --------- ---------/--------  -----
F-0190 HGA_AZ_ANG         0     0  0.00E+00  0.00E+00  0.00E+00/0.00E+00      4
F-0195 HGA_EL_ANG         0     0  0.00E+00  0.00E+00  0.00E+00/0.00E+00      4
L-0034 MOT1_RANGING       0     0                                            54
L-0036 MOT1_TLM_MOD       0     0                                            54
L-0037 MOT1_USO_ENA       0     0                                            54
L-0044 MOT2_RANGING       0     0                                            54
L-0046 MOT2_TLM_MOD       1     1                                            54
L-0047 MOT2_USO_ENA       1     1                                            54
L-0053 RF_SW_OSC          1     1                                            54
L-0060 TWTA1_FILMNT       1     1                                            54
L-0061 TWTA1_HV           1     1                                            54
L-0070 TWTA2_FILMNT       1     1                                            54
L-0071 TWTA2_HV           0     0                                            54
L-0080 USO_PWR            0     0                                            54
L-0090 KaBLE_STAT         1     1                                            54
L-0091 KaBLE_ENABLE       0     0                                            54
L-0111 MOT1_RCV_AGC     255   255 -1.53E+02 -1.53E+02 -1.53E+02/0.00E+00     27
L-0121 MOT2_RCV_AGC     255   255 -1.53E+02 -1.53E+02 -1.53E+02/0.00E+00     27
L-0131 TWTA1_HLX_I        0     0 -3.84E-02 -3.84E-02 -3.84E-02/0.00E+00     27
L-0132 TWTA1_ANOD_V     243   245 -5.19E+02 -5.15E+02 -5.17E+02/2.33E-01     27
L-0141 TWTA2_HLX_I       39    42  7.25E-01  7.85E-01  7.64E-01/2.66E-03     27
L-0142 TWTA2_ANOD_V      63    65 -8.62E+02 -8.58E+02 -8.61E+02/2.73E-01     27
L-0200 USO_REG_V        203   205  4.06E+00  4.10E+00  4.08E+00/1.83E-03     27
L-0201 USO_OVEN_V       122   124  2.44E+00  2.48E+00  2.46E+00/1.28E-03     27
T-0302 EPC1_T           126   128  2.42E+01  2.50E+01  2.45E+01/1.46E-01      5
T-0303 EPC2_T           113   115  2.98E+01  3.07E+01  3.03E+01/1.73E-01      5
T-0310 RF_ISOLTR1_T     184   188  2.53E+01  2.68E+01  2.60E+01/6.94E-02     27
T-0311 RF_ISOLTR2_T     179   182  2.75E+01  2.86E+01  2.80E+01/7.26E-02     27
T-0312 TWT1_T           184   188  2.53E+01  2.68E+01  2.61E+01/2.77E-01      5
T-0313 TWT2_T           171   175  3.00E+01  3.14E+01  3.07E+01/2.28E-01      5
T-0315 USO_T            185   187  7.63E+00  8.30E+00  7.90E+00/1.12E-01      5
T-0316 KA_AMP_T         158   163  3.42E+01  3.60E+01  3.47E+01/1.27E-01     13

   L-0200 and L-0201 are nominal. T-0315 continues to drop, but slowly.


Open Loop Data
==============

  Data Collection
  ---------------

      No new data.

  Analysis
  --------

     Joe Twicken has completed processing of all occultation data.  In the
past couple days he has processed ODRs from occultations 119-142 (both
ingress and egress) using the latest GDN files from Goddard.  He reports
two problems:

Occultation #136i (start time 1998/105T01:05:20, DSS 45) suffers from the
sawtooth problem. The egress ODR could not be processed to completion due to
low signal level. The trajectory for this orbit also does not seem very good
(earlier analysis of data from 1998/104, using the same GDN reconstruction,
had been judged less satisfactory than others).

It appears that there were problems with a complete spectrum full of frequency
harmonics during the ingress occultations #139 (start time 1998/106T12:02:20)
and #141 (1998/107T11:21:32) at DSS 65.  In the first case, there is a very
well defined set of harmonics with what looks to be 20 Hz spacing. In the
second case, the harmonics are not as well defined.  In both cases, however,
the spurs can even be seen pretty clearly in the PREPLOOK spectra.


Closed Loop Data
================

  Data Collection
  ---------------

     We have obtained the following files from OSCAR:

                Original File Name               Name on SOPC
                -------------------------------- ------------
       ODF:     98126O127.DAT;1                  8125126A.ODF
       TDF:     98118T119.DAT;1                  8118119A.TDF
       TDF:     98119T120.DAT;1                  8119120A.TDF
       TDF:     98120T121.DAT;1                  8120121A.TDF
       TDF:     98121T122.DAT;1                  8121122A.TDF
       TDF:     98122T123.DAT;1                  8122123A.TDF
       TDF:     98123T125.DAT;1                  8123125A.TDF


Ancillary Data
==============

  Data Collection
  ---------------

      No new data.